what is the molecular weight of a gas if it weights 0.7g and occupies a volume of 1 litre?
1
Expert's answer
2012-12-04T06:23:47-0500
m V
Gas------------>Gas
Mw 22.4 (const)

Mw = m*22.4/V

Mw = 0.7*22.4/1 = 15,68